OpenStudy (anonymous):

Three consequtive terms of an arithmetic sequence have a sum of 12 and a product of -80. Find the terms

OpenStudy (anonymous):

middle term a, difference d a-d+a+a+d=12 a=4 (4-d)4(4+d)=-80 16-d^2=-20 d^2=36 d=6 or d=-6 Terms: -2, 4, 10 or 10, 4, -2

OpenStudy (anonymous):

-2,4,10